# Geometric properties of the additional third-order transitions in the two-dimensional Potts model

### v2

@@ Abstract
keywords:
potts model , third-order transition , Swendsen-Wang Algorithm
[…]
[label1]organization=College of Sciences, addressline=Xi’an University of Science and Technology, city=Xi’an, postcode=710054, country=China
[…]
[label2]organization=School of Systems Science, addressline=Beijing Normal University, city=Beijing, postcode=100875, country=China
[…]
[label3]organization=International Academic Center of Complex Systems, addressline=Beijing Normal University, city=ZhuHai, postcode=519087, country=China
[…]
[label4]organization=Department of Systems Science, Faculty of Arts and Sciences, addressline=Beijing Normal University, city=ZhuHai, postcode=519087, country=China
[…]
[label5]organization=2020 X-Lab, Shanghai Institute of Microsystem and Information Technology, addressline=Chinese Academy of Sciences, city=ShangHai, country=China
1 Introduction
@@
The Potts model [1, 2], a generalization of the Ising model [3, 4], has long been fundamental in statistical mechanics for studying phase transitions and critical phenomena in complex systems. Originally developed to explore magnetic properties [2], the Potts model has found applications across diverse fields, including optimizing communication networks like Wireless Body Area Networks and MESH networks [5]. Recent studies have expanded its scope to disciplines such as biophysics and statistical physics. For instance, Bae and Tai [6] introduced a four-phase Potts model for image segmentation, while Morcos et al. [7] applied it to protein structure prediction using direct coupling analysis. The model’s relevance has recently been extended to machine learning, where Rende and colleagues [8] mapped the self-attention mechanism onto a generalized Potts model, demonstrating its effectiveness in solving the inverse Potts problem. These developments highlight the model’s expanding interdisciplinary impact.
@@
Traditional statistical physics and thermodynamics have been highly successful in explaining the behavior of phase transitions in systems. A phase transition is a cooperative phenomenon involving multiple factors such as temperature and pressure. The structure and properties of a system evolve in response to variations in certain order parameters, primarily temperature[9]. However, in the $q$-state Potts model ($q\geq 3$), when $q=2$, the model reduces to the Ising model[2]. For $1\leq q\leq 4$, the phase transition of the Potts model is continuous[10, 11], while for $q>4$, the phase transition becomes discontinuous[11, 12, 13, 14, 15, 16, 17]. By incorporating bond percolation theory[18, 19, 20, 21], we can easily determine whether a phase transition is continuous or not. Percolation theory [22, 23, 24, 25] is the natural framework to study the properties of cluster-like structures of a system. Extensive research on percolation, especially focusing on both bond and site percolation in two and three dimensions, has been conducted, with significant contributions from Youjin Deng and his collaborators, covering critical transitions in percolation models [26, 27, 28, 29, 30]. As we known, the early warning of phase transitions hold theoretical significance and practical value[31, 32]. It raises the question of whether there exist abnormal behaviors that can act as precursors to the main phase transition.
@@
Over the past few decades, microcanonical analysis has emerged as a valuable approach for identifying phase transitions in physical systems [33, 34]. Microcanonical inflection-point analysis (MIPA) has been recognized as an effective method for studying phase transitions in finite-size systems. Qi and Bachmann [35] expanded this approach to effectively identify higher-order transitions, distinguishing between independent and dependent transitions. The inflection point in a higher-order derivative of the microcanonical entropy specifically marks the point at which the monotonicity of the function changes, signaling the occurrence of a phase transition. Independent transitions, akin to conventional phase transitions, occur independently of other cooperative processes within the system. Dependent transitions, in contrast, are contingent upon the occurrence of lower-order transitions and take place at higher energy levels, representing higher-order phenomena.
@@
Utilizing the exact density of states (DOS) of the Ising model [36], Sitarachu performed a comprehensive analysis of 1D and 2D finite-size Ising models, identifying higher-order transitions in the 2D system [37]. By using geometric order parameters in the canonical ensemble, Sitarachu et al. effectively demonstrated the correspondence between canonical and microcanonical ensemble methods for identifying third-order transitions [38]. Their analysis identified two forms of third-order phase transitions: independent and dependent transitions. The third-order independent transition occurs before the critical phase transition and is geometrically associated with the peak in the number of isolated spins, which disrupts the order of the system. This maximum is observed below the critical temperature. In contrast, the third-order dependent transition emerges beyond the critical point and is associated with a local minimum in the first derivative of the average cluster size with respect to temperature, indicating a shift in the decay rate of clusters in the disordered phase. This dependent transition is marked by a minimum above the critical temperature.
@@
This paper is organized as follows: Section 2 introduces the Potts model and the Swendsen-Wang Algorithm, along with a detailed discussion on the third-order dependent phase transition observed in the system. Section 3 explores phase transitions, with Section 3.1 examining critical transitions—second-order for $q=3$ and $q=4$, and first-order for $q=6$ and $q=8$. Section 3.2 examines the third-order independent transition and the definition of isolated spins in the Potts model, while Section 3.3 details the use of the average perimeter as the order parameter for the third-order dependent transition. Section 4 concludes by discussing the results.
2 The model and method
2.1 Potts Model
The $q$-state Potts model [39], a generalization of the Ising model, is extensively utilized to study phase transitions in statistical mechanics. In this model, spins on a lattice may assume one of $q$ discrete states, as opposed to the binary states in the Ising model. There is no external magnetic field in this model. The energy for a specific configuration of spins $X=(s_{1},s_{2},\dots,s_{N})$, on a lattice with $N=L\times L$ sites, is expressed as:
@@ 2.1 Potts Model
where $s_{i}$ represents the spin at site $i$, which may take values from 00 to $q-1$, and $\delta(s_{i},s_{j})$ is the Kronecker delta function, which equals 1 when $s_{i}=s_{j}$ and 0 otherwise. The summation is performed over all nearest-neighbor pairs $\langle i,j\rangle$, and $J>0$ is the coupling constant that favors alignment of neighboring spins, reflecting ferromagnetic interactions. When $q=2$, the Potts model reduces to the Ising model, while larger values of $q$ lead to increasingly complex behavior and phase transition dynamics. The model has been thoroughly investigated for its applications in understanding critical phenomena and phase transitions [11].
2.2 Swendsen-Wang Algorithm
It is well established [18] that the Potts model is intricately connected to problems of connectivity and percolation in graph theory. The Fortuin-Kasteleyn transformation [40] enables a mapping of the original model, which suffers from slow critical relaxation, into one where such relaxation effects are significantly mitigated.The transformation replaces each pair of interacting Potts spins on the lattice with a bond on an equivalent lattice with the probability:
@@ 2.2 Swendsen-Wang Algorithm
where $K={J}/{k_{B}T}$, and $J=1$. A bond is placed with probability $p$ only when $\sigma_{i}=\sigma_{j}$, indicating that the spins of the nearest neighbors are in the same state.This implies that bonds are only formed with a non-zero probability when the corresponding pair of spins on the original lattice are in the same state. The process must be repeated for all pairs of spins, resulting in a lattice where bonds connect certain sites, forming clusters of varying sizes and shapes.
@@
The first use of the Fortuin–Kasteleyn transformation in Monte Carlo simulations was made by Swendsen and Wang [41]. The process involves traversing the lattice and placing bonds between each pair of spins with the probability given by equation (2). The Hoshen–Kopelman [42] method is then employed to identify all clusters of sites connected by the bond network. Each cluster is subsequently assigned a new spin value from the set $q=\{0,1,2,\dots,q-1\}$, ensuring that all sites within a given cluster share the same randomly chosen spin value. This constitutes one Monte Carlo step in our algorithm. In the next Monte Carlo step, based on the spin configuration from the previous step, new bonds are re-established between neighboring spins according to the predefined probability. The Hoshen-Kopelman algorithm[42] is once again employed to identify clusters, after each cluster is reassigned a new spin value uniformly from the set $q=\{0,1,2,\dots,q-1\}$. This process repeats iteratively until the desired number of Monte Carlo steps is reached.It is important to note that in our algorithm, a dedicated array is used to store the bond states between spins. This array is reset at the start of each Monte Carlo step, ensuring that the bonding configuration is entirely refreshed, with no correlation to the previous step’s bonds.
@@
In our algorithm, we emphasize that during each Monte Carlo (MC) step, the Hoshen-Kopelman (H-K) algorithm is applied twice for cluster statistics: first to identify clusters within the system, followed by a spin-flipping process in which F-K clusters are defined by introducing a bond between two nearest-neighbor spins with identical orientations, based on the probability outlined in equation (2). However, once the F-K clusters are established, it is essential to ascertain the true physical state of the system to compute the order parameters accurately. This requires the formation of clusters solely from nearest-neighbor spins that share the same orientation after the spin-flipping operations, ensuring that the final cluster information aligns with the actual physical system. Thus, while the construction of F-K clusters addresses the issue of critical slowing down during spin flips, the accurate computation of physical quantities in the system necessitates considering the actual physical clusters present.
@@
Since the probability of placing a bond between pairs of sites depends on temperature, it is clear that the resultant cluster distributions will vary dramatically with temperature. At very high temperatures, the clusters tend to be quite small. At very low temperatures, virtually all sites with nearest neighbors in the same state will end up in the same cluster, leading to oscillations between similar structures. Near the critical point, a diverse array of clusters emerges, resulting in each configuration differing significantly from its predecessor. This effectively reduces the phenomenon of critical slowing down. By employing the Swendsen-Wang (S-W) algorithm, the critical slowing down is significantly mitigated, facilitating faster convergence and more efficient sampling in Monte Carlo simulations[43].
@@
If we aim to study the order of phase transitions, the F-K transformation is an effective tool for this purpose. In the F-K clusters, the probability of a randomly selected spin belonging to the largest cluster, denoted as $\langle P_{\infty}\rangle=\langle n_{\infty}\rangle/{L^{2}}$, where $\langle n_{\infty}\rangle$ represents the size of the largest cluster, shows behavior that is consistent with the magnetization $\langle M\rangle$ as calculated from real clusters when $T\leq T_{c}$ [44]. However, in the paramagnetic phase ($T>T_{c}$), discrepancies arise between these two quantities, as many clusters contribute to the magnetization $\langle M\rangle$, rather than solely the largest cluster, which is represented by $\langle P_{\infty}\rangle$. Fluctuation quantities, such as specific heat and susceptibility, exhibit related differences that arise from distinct contributions of clusters. In particular, one must separate the contributions from clusters smaller than the largest one and those from the size of the largest cluster itself[45]. To identify additional signals of higher-order phase transitions, especially in the paramagnetic phase, it is crucial to seek an alternative and appropriate order parameter that can more effectively capture subtle changes associated with these transitions[38].
2.3 Third-order transitions
@@
Microcanonical inflection-point analysis demonstrates the occurrence of a third-order independent transition before the critical phase transition. Geometrically, this third-order independent transition corresponds to a peak in the count of isolated spins, observed just before the phase transition. In the Ising model, the third-order transition is observed at $T_{\text{in}}\approx 2.229$ [38]. However, when applying the same definition of isolated spins from the Ising model to the Potts model, the number of isolated spins increases monotonically with temperature, without showing evidence of a third-order independent transition. This is because, in the $q$-state Potts model (for $q\geq 3$), spins can adopt more than two states. If isolated spins are defined solely based on having a different orientation from their four nearest neighbors, as depicted in Fig. 1 (a), their count will continue increasing monotonically with temperature, without showing any anomalous behavior indicative of a third-order independent transition.
@@
Therefore, in the $q$-state Potts model (for $q\geq 3$), it is necessary to redefine the concept of isolated spins. In the Ising model, it is observed that, due to the nature of the system, an isolated spin is surrounded by four nearest neighbors that have the same orientation, which is opposite to that of the spin itself. This leads to the hypothesis that the role of an isolated spin is to serve as a disrupter of order within the system. Extending this logic to the $q$-state Potts model, isolated spins are similarly defined. In an ordered cluster, an isolated spin serves as a disrupter, with a different orientation from its four nearest neighbors, as shown in Fig. 1 (b), all of which belong to the same cluster. The isolated spin thus behaves like a "nail" driven into an ordered cluster, demonstrating its function as a source of disorder within an otherwise ordered system.
@@
Figure 2: Cluster perimeter illustration for the $q=3$ Potts Model. This figure demonstrates the relationship between cluster size and perimeter in the $q=3$ Potts model. While the cluster size only indicates the number of spins within a cluster, the perimeter more accurately captures the boundary characteristics of the cluster. Clusters with the same number of spins can exhibit different configurations, resulting in distinct perimeters. This distinction makes the perimeter a more precise and effective metric for analyzing cluster configurations and identifying signals of higher-order phase transitions.
@@
Microcanonical inflection-point analysis demonstrates that an additional third-order dependent transition occurs after the critical phase transition, occurring specifically within the paramagnetic phase, at a temperature distinct from the critical transition temperature. In the canonical ensemble, this corresponds to a slight shift in the rate at which the average cluster size decreases after the critical phase transition in the Ising model, specifically when $T>T_{c}$. Mathematically, this is identified by a local minimum in the first derivative of the average cluster size with respect to temperature. The temperature at which this minimum occurs indicates the occurrence of the third-order dependent transition, which, in the Ising model, is found to be $T_{\text{de}}\approx 2.567$[38]. Following this logic, this approach is extended to the $q$-state Potts model (for $q\geq 3$) to search for the signal of a third-order dependent transition.
@@
Subsequently, several order parameters were investigated in the $q$-state Potts model (for $q\geq 3$), and ultimately, the average perimeter size of clusters was identified as a key indicator for detecting the third-order dependent transition in the model. The process of identifying the average perimeter as an order parameter was somewhat unexpected. After realizing that the average cluster size could not pinpoint the third-order dependent transition in the $q$-state Potts model, an analysis of this parameter was conducted in detail. The average cluster size represents the average area of clusters, indicating the number of spins within each cluster. Since area failed to capture the signal, the focus was shifted to examining the average perimeter, which characterizes the number of spins on the boundary of each cluster,as shown in Fig. 2. Interestingly, during the analysis of the temperature dependence of the average perimeter, a similar trend was observed to that of the average cluster size in the 2D Ising model.

### v3

@@ Abstract
I Introduction
@@
The Potts model Potts (1952); Wu (1984) generalizes the Ising model to multiple states Onsager (1944); Brush (1967), with different spin states represented by integer values of $q$. It has become an important framework in statistical mechanics, contributing significantly to the understanding of phase transitions and critical phenomena in complex systems. Originally developed to explore magnetic properties Wu (1984), the Potts model has found applications across diverse fields, including optimizing communication networks like Wireless Body Area Networks and MESH networks Paszkiewicz and Węgrzyn (2020). Recent studies have expanded its scope to disciplines such as biophysics and statistical physics. For instance, Bae and Tai Tai et al. (2023) introduced a four-phase Potts model for image segmentation, while Morcos et al. Morcos et al. (2011) applied it to protein structure prediction using direct coupling analysis. The model’s relevance has recently been extended to machine learning, where Rende and colleagues Rende et al. (2024) mapped the self-attention mechanism onto a generalized Potts model, demonstrating its effectiveness in solving the inverse Potts problem. These developments highlight the model’s expanding interdisciplinary impact.
@@
Traditional statistical physics and thermodynamics have been highly successful in explaining the behavior of phase transitions in systems. A phase transition is a cooperative phenomenon involving multiple factors such as temperature and pressure. The structure and properties of a system evolve in response to variations in certain order parameters, primarily temperature Pathria (2001). In the $q$-state Potts model, for $1\leq q\leq 4$, the phase transition of the Potts model is continuous Duminil-Copin et al. (2017); Baxter (1973), while for $q>4$, the phase transition becomes discontinuous Baxter (1973); Duminil-Copin et al. (2021); Iglói and Carlon (1999); Hamer (1981); Iglói and Sólyom (1983); Loureiro et al. (2010, 2012). By incorporating bond percolation theory Fortuin and Kasteleyn (1972); Stephen (1976); Wu (1978); Lee et al. (2005), we can easily determine whether a phase transition is continuous or not. Percolation theory Stauffer and Aharony (2018); Grimmett (1999); Essam (1980); Saberi (2015) is the natural framework to study the properties of cluster-like structures of a system. Extensive research on percolation, especially focusing on both bond and site percolation in two and three dimensions, has been conducted, with significant contributions from Youjin Deng and his collaborators, uncovering phase transitions in percolation models Wang et al. (2013); Feng et al. (2008); Deng and Blöte (2003, 2005); Xu et al. (2014). As we know, the early warning of phase transitions hold theoretical significance and practical value Scheffer et al. (2009); Freeman and CG (2012). It raises the question of whether there exist typical and universal behaviors that can act as precursors to main phase transitions, and whether these have the potential to be applied beyond classical magnetic systems, such as artificial swarms Bechinger et al. (2016) or active bacterial colonies Copeland and Weibel (2009); Qi et al. (2022).
@@
Over the past few decades, microcanonical analysis has emerged as a valuable approach for identifying phase transitions in physical systems Gross (2001); Junghans et al. (2006). Microcanonical inflection-point analysis (MIPA) has been recognized as an effective method for studying phase transitions in finite-size systems. Qi and Bachmann Qi and Bachmann (2018) expanded this approach to effectively identify higher-order transitions, distinguishing between independent and dependent transitions. The inflection point in a higher-order derivative of the microcanonical entropy specifically marks the point at which the monotonicity of the function changes, signaling the occurrence of a higher-order transition. Independent transitions, akin to conventional phase transitions, occur independently of other cooperative processes within the system. Dependent transitions, in contrast, are contingent upon the occurrence of lower-order transitions and take place at higher energy levels, representing higher-order phenomena.
@@
Utilizing the exact density of states (DOS) of the Ising model Beale (1996), Sitarachu performed a comprehensive analysis of 1D and 2D finite-size Ising models, identifying higher-order transitions in the 2D system Sitarachu et al. (2020). By using geometric order parameters in the canonical ensemble, Sitarachu et al. effectively demonstrated the correspondence between canonical and microcanonical ensemble methods for identifying third-order transitions Sitarachu and Bachmann (2022). Their analysis identified two forms of third-order transitions: independent and dependent transitions, and determined how to use geometric order parameters to locate the positions of these transitions. The third-order independent transition occurs before the phase transition and is geometrically associated with the peak in the number of isolated spins, which disrupts the order of the system. This maximum is observed below the transition temperature. In contrast, the third-order dependent transition emerges beyond the transition point and is associated with a local minimum in the first derivative of the average cluster size with respect to temperature, indicating a shift in the decay rate of clusters in the disordered phase. This dependent transition is marked by a minimum above the transition temperature.
@@
This paper is organized as follows: Section II provides an overview of the Potts model and the Swendsen-Wang algorithm, along with an analysis of the third-order transitions identified using geometric order parameters. Section III examines the main phase transition order for different values of $q$ and investigates the presence of higher-order transitions, emphasizing their dependence on the phase transition type. Finally, Section IV summarizes the article.
II The model and method
II.1 Potts Model
The $q$-state Potts model Wu (1982), a generalization of the Ising model, is extensively utilized to study phase transitions in statistical mechanics. In this model, spins on a lattice may assume one of $q$ discrete states, as opposed to the binary states in the Ising model. There is no external magnetic field in this model. The energy for a specific configuration of spins $X=(s_{1},s_{2},\dots,s_{N})$, on a lattice with $N=L\times L$ sites, is expressed as:
@@ II.1 Potts Model
where $s_{i}$ represents the spin at site $i$, which may take integer values from 00 to $q-1$, and $\delta(s_{i},s_{j})$ is the Kronecker delta function, which equals $1$ when $s_{i}=s_{j}$ and 00 otherwise. The summation is performed over all nearest-neighbor pairs $\langle i,j\rangle$, and $J>0$ is the coupling constant that favors alignment of neighboring spins, reflecting ferromagnetic interactions. When $q=2$, the Potts model reduces to the Ising model, while larger values of $q$ lead to increasingly complex behavior and phase transition dynamics. The model has been thoroughly investigated for its applications in understanding critical phenomena and phase transitions Baxter (1973).
II.2 Swendsen-Wang Algorithm
It is well established Fortuin and Kasteleyn (1972) that the Potts model is intricately connected to problems of connectivity and percolation in graph theory. The Fortuin-Kasteleyn transformation Kasteleyn and Fortuin (1969) enables a mapping of the original model, which suffers from critical slowdown, into one where such slowdown effects are significantly mitigated. The transformation places a bond between each pair of interacting Potts spins on the lattice with the probability:
@@ II.2 Swendsen-Wang Algorithm
where $K={J}/{k_{B}T}$, and $J=1$. A bond is placed with probability $p$ only when $\sigma_{i}=\sigma_{j}$, indicating that the spins of the nearest neighbors are in the same state. This implies that bonds are only formed with a non-zero only when the probability when the corresponding pair of spins on the original lattice are in the same state. The process must be repeated for all pairs of spins, resulting in a lattice where bonds connect certain sites, forming clusters of varying sizes and shapes.
@@
Swendsen and Wang Swendsen and Wang (1987) successfully applied the Fortuin–Kasteleyn transformation in Monte Carlo simulations. The process involves traversing the lattice and placing bonds between each pair of spins with the probability given by Equation (2). The Hoshen–Kopelman Hoshen and Kopelman (1976) method is then used to identify all clusters of sites connected by the bond network. Subsequently, each cluster is assigned a new spin value from the set $q=\{0,1,2,\dots,q-1\}$, ensuring that all sites within a given cluster share the same randomly chosen spin value. This constitutes one Monte Carlo step in our algorithm. In the next Monte Carlo step, based on the spin configuration from the previous step, new bonds are re-established between neighboring spins according to the predefined probability. The Hoshen-Kopelman algorithm Hoshen and Kopelman (1976) is once again used to identify clusters, after each cluster is reassigned a new spin value uniformly from the set $q=\{0,1,2,\dots,q-1\}$. This process repeats iteratively until the desired number of Monte Carlo steps is reached. It is important to note that in our algorithm, a dedicated array is used to store the bond states between spins. This array is reset at the start of each Monte Carlo step, ensuring that the bonding configuration is entirely refreshed, with no correlation to the bonds in the previous step.
@@
In our algorithm, we emphasize that during each Monte Carlo (MC) step, the Hoshen-Kopelman (H-K) algorithm is applied twice for cluster statistics: first to identify clusters within the system, followed by a spin-flipping process in which F-K clusters are defined by introducing a bond between two nearest-neighbor spins with identical orientations, based on the probability outlined in equation (2). However, once the F-K clusters are established, it is essential to ascertain the true physical state of the system to compute the order parameters accurately. This requires the formation of clusters solely from nearest-neighbor spins that share the same orientation after the spin-flipping operations, ensuring that the final cluster information aligns with the actual physical system. Thus, while the construction of F-K clusters addresses the issue of critical slowdown during spin flips, the accurate computation of physical quantities in the system necessitates considering the actual physical clusters present.
@@
Since the probability of placing a bond between pairs of sites depends on temperature, it is clear that the resultant cluster distributions will vary dramatically with temperature. At very high temperatures, the clusters tend to be quite small. At very low temperatures, virtually all sites with closest neighbors in the same state will end up in the same cluster, leading to oscillations between similar structures. Near the transition point, a diverse array of clusters emerges, resulting in each configuration differing significantly from its predecessor. This effectively reduces the phenomenon of critical slowdown. Using the Swendsen-Wang algorithm (S-W), critical slowdown is significantly mitigated, facilitating faster convergence and more efficient sampling in Monte Carlo simulations Newman and Barkema (1999).
@@
If we aim to study the behavior of phase transitions, the F-K transformation is an effective tool for this purpose. In the F-K clusters, the probability of a randomly selected spin belonging to the largest cluster, denoted as $\langle P_{\infty}\rangle=\langle n_{\infty}\rangle/L^{2}$, where $\langle n_{\infty}\rangle$ represents the size of the largest cluster, exhibits behavior consistent with the magnetization $\langle M\rangle$ calculated from real clusters for temperatures $T\leq T_{\text{c}}$. However, for $T>T_{\text{c}}$, the value of $\langle P_{\infty}\rangle$ is slightly smaller than that of $\langle M\rangle$ De Meo et al. (1990). In the paramagnetic phase ($T>T_{\text{c}}$), discrepancies arise between these two quantities, as many clusters contribute to the magnetization $\langle M\rangle$, rather than solely the largest cluster, which is represented by $\langle P_{\infty}\rangle$. Fluctuation quantities, such as specific heat and susceptibility, exhibit related differences that arise from distinct contributions of clusters. In particular, contributions must be separated from clusters smaller than the largest and those from the size of the largest cluster itself Landau and Binder (2021). To identify additional signals of higher-order phase transitions, especially in the paramagnetic phase, it is crucial to seek an alternative and appropriate order parameter that can more effectively capture subtle changes associated with these transitions Sitarachu and Bachmann (2022).
II.3 Third-order transitions
@@
Microcanonical inflection-point analysis demonstrates the occurrence of a third-order independent transition prior to the phase transition temperature $T_{\text{c}}$. Geometrically, this third-order independent transition corresponds to a peak in the count of isolated spins, observed at a temperature $T_{\text{ind}}$ lower than the phase transition temperature $T_{\text{c}}$. In the Ising model, the third-order transition is observed at $T_{\text{ind}}\approx 2.229$ Sitarachu and Bachmann (2022). However, when the same definition of isolated spins from the Ising model is applied to the Potts model, the number of isolated spins increases monotonically with temperature, without showing evidence of a third-order independent transition. We speculate that this may be because, in the Potts model in the $q$ state (for $q\geq 3$), the spins can adopt more than two states. If isolated spins are simply defined as those with a spin direction different from that of the four nearest neighbors, then for the $q$-state Potts model with $q\geq 3$, the system will contain two types of single-spin clusters: one as shown in Fig. 1 (a) and the other as shown in Fig. 1 (b). For the type of single-spin shown in Fig. 1 (a), the ordered state of the region it resides in has already been broken. Therefore, it is meaningless to include it in the statistics. We only need to consider the isolated spins as shown in Fig. 1 (b).
@@
As a result, in the $q$-state Potts model (for $q\geq 3$), it is necessary to redefine the concept of isolated spins. In the Ising model, it is observed that, due to the two-state nature of the Ising model, an isolated spin is surrounded by four nearest neighbors that have the same orientation, which is opposite to that of the spin itself. This leads to the hypothesis that the role of an isolated spin is to serve as a disrupter of order within the system. Extending this logic to the $q$-state Potts model, isolated spins are similarly defined. In an ordered cluster, an isolated spin serves as a disrupter, with a different orientation from its four nearest neighbors, as shown in Fig. 1 (b), all of which belong to the same cluster. The isolated spin thus behaves like a "nail" driven into an ordered cluster, demonstrating its function as a source of disorder within an otherwise ordered system.
@@
Figure 2: Cluster perimeter illustration for the $q=3$ Potts model. This figure demonstrates the relationship between cluster size and perimeter in the $q=3$ Potts model. While the cluster size only indicates the number of spins within a cluster, the perimeter more accurately captures the boundary characteristics of the cluster. Clusters with the same number of spins can exhibit different configurations, resulting in distinct perimeters. This distinction makes the perimeter a more precise and effective metric for analyzing cluster configurations and identifying signals of higher-order transitions.
@@
Microcanonical inflection-point analysis demonstrates that an additional third-order dependent transition occurs within the paramagnetic phase, at a temperature $T_{\text{de}}$ distinct from the phase transition temperature $T_{\text{c}}$, where $T_{\text{de}}$ is larger than $T_{\text{c}}$. In the canonical ensemble, this corresponds to a slight shift in the rate at which the average cluster size decreases after the critical phase transition in the Ising model, specifically when $T>T_{\text{c}}$. Mathematically, this is identified by a local minimum in the first derivative of the average cluster size with respect to temperature. The temperature at which this minimum occurs indicates the occurrence of the third-order dependent transition, which, in the Ising model, is found to be $T_{\text{de}}\approx 2.567$ Sitarachu and Bachmann (2022). Following this logic, this approach is extended to the $q$-state Potts model (for $q\geq 3$) to search for the signal of a third-order dependent transition, but the attempt did not yield any successful results.
@@
Subsequently, several order parameters were investigated in the $q$-state Potts model (for $q\geq 3$), and ultimately, the average perimeter of clusters was identified as a indicator for detecting the third-order dependent transition in the model. The process of identifying the average perimeter as an order parameter was somewhat unexpected. After realizing that the average cluster size could not pinpoint the third-order dependent transition in the $q$-state Potts model, an analysis of this parameter was conducted in detail. The average cluster size represents the average area of clusters, indicating the number of spins within each cluster. Since area failed to capture the signal, the focus was shifted to examining the average perimeter, which characterizes the number of spins on the boundary of each cluster,as shown in Fig. 2. Interestingly, during the analysis of the temperature dependence of the average perimeter, a similar trend was observed to that of the average cluster size in the 2D Ising model.